### v2

@@ Abstract
Keywords:
model order reduction parametric dynamical systems Low-rank tensor completion tensor decomposition
MSC:
65P99 15A69
@@ 1 Introduction
we introduce a Cartesian grid on $\mathcal{A}$ by distributing $K_{i}$ nodes $\{\widehat{\alpha}_{i}^{j}\}_{j=1,\dots,K_{i}}$ within each of the intervals $[\alpha_{i}^{\min},\alpha_{i}^{\max}]$ for $i=1,\dots,D$. The nodes of the grid form the set
@@
The discrete parametric solution manifold is defined by the set of trajectories for all parameters $\widehat{\boldsymbol{\alpha}}\in{\widehat{\mathcal{A}}}$, where each trajectory is given by the snapshot matrix $\Phi(\widehat{\boldsymbol{\alpha}})$. Alternatively, this manifold can be represented by the multi-dimensional array
@@
It is clear that for a high dimension $D$ or a fine parameter grid ${\widehat{\mathcal{A}}}$, working directly with $\boldsymbol{\Phi}$ is prohibitively expensive, if possible at all. Therefore, one approach to model order reduction for a system like (1) is based on the assumption that $\boldsymbol{\Phi}$ can be effectively approximated by a low-rank tensor $\widetilde{\boldsymbol{\Phi}}$ in one of the commonly used formats. Depending on the definition of tensor rank, various tensor decomposition formats can be employed, including CANDECOMP/PARAFAC (CP), Tucker (HOSVD), or Tensor Train (TT). While known rigorous analyses revealing how tensor ranks depend on the properties of a parameterized differential equation and the targeted accuracy of the solution manifold recovery, or its smoothness, are very limited (see, e.g., bachmayr2017kolmogorov and discussions in khoromskij2011tensor ; nouy2017low ; bachmayr2023low ), numerical evidence suggests that for many practical parametric PDEs, solutions to such problems or statistics derived from them are well approximated in low-rank formats kressner2011low ; ballani2015hierarchical ; dolgov2015polynomial ; ballani2016reduced ; eigel2017adaptive ; bachmayr2018parametric ; dolgov2019hybrid ; glau2020low ; mamonov2022interpolatory ; mamonov2024tensorial .
@@
with a prescribed $\varepsilon\geq 0$. The inexact completion (7) is a commonly used problem setup for the case of noisy data candes2010matrix and this is the formulation we are interested here. For CP, Tucker, and TT tensor ranks, the completion problems (6) and (7) are NP-hard. Building on the success of solving low-rank matrix completion problems candes2010power ; cai2010singular ; candes2012exact ; vandereycken2013low ; jain2013low ; nguyen2019low , a popular approach is to relax (6)–(7) into convex optimization problems signoretto2010nuclear ; gandy2011tensor ; bengua2017efficient . Other approaches to tensor completion include ALS methods yokota2016smooth ; grasedyck2019stable , Riemannian optimization kressner2014low ; steinlechner2016riemannian , Bayesian methods zhao2015bayesian ; zhao2015bayesian2 , and projection methods rauhut2015tensor , with applications ranging from video recovery to seismic data reconstruction bengua2017efficient ; liu2022efficient .
@@
as a slice, hence the name slice sampling. For the problem (1), we have $C=2$, and the slice (8) coincides with the snapshot matrix (2), where $\boldsymbol{\alpha}=\big{(}\widehat{\alpha}_{1}^{k_{1}},\dots,\widehat{\alpha}_{D}^{k_{D}}\big{)}^{T}$. The sampled indices $k_{1},\dots,k_{D}$ form a sparse subset of the grid ${\widehat{\mathcal{A}}}$.
@@
The reconstruction of scalar output quantities of parametric solutions in tensor form from incomplete observations was addressed in ballani2015hierarchical ; glau2020low . In ballani2015hierarchical the authors employed a tensor cross approximation, while glau2020low applied TT-completion via Riemannian optimization to recover an option pricing statistic from solutions of parametrized Heston and multi-dimensional Black-Scholes models. Additionally, a comparison of TT-cross interpolation and TT-completion for a parameterized diffusion equation in steinlechner2016riemannian demonstrated that TT-completion requires fewer PDE solver executions to find a low-rank approximation of a particular solution functional.
@@ 2 Slice sampling tensor completion
$\mathcal{D}=\left\{\Phi_{\mathbf{i},\mathbf{j}}=\Phi_{i_{1},\ldots,i_{C},j_{1},\ldots,j_{D}}\;\left|\;\forall\,\mathbf{i}\in\Omega_{C},\;\forall\,\mathbf{j}\in\widetilde{\Omega}_{D}\right.\right\},$ (12)
@@ 2.2 Tensor-train completion for partially sampled modes
$[\boldsymbol{\Psi}\times_{n}\mathrm{M}]_{i_{1},\ldots,i_{n-1},j,i_{n+1},\ldots,i_{d}}=\sum_{k=1}^{m_{n}}\boldsymbol{\Psi}_{i_{1},\ldots,i_{n-1},k,i_{n+1},\ldots,i_{d}}\mathrm{M}_{jk},\quad j=1,\ldots,m.$ (19)
@@
$\mathcal{D}^{q}=\left\{\boldsymbol{\Phi}^{q}_{\mathbf{i},\mathbf{j}}\;\left|\;\forall\mathbf{i}\in\Omega_{C}^{q},\;\forall\mathbf{j}\in\widetilde{\Omega}_{D}\right.\right\},$ (21)
@@ 2.5 Accuracy of HTT completion
$\boldsymbol{\Phi}-\widehat{\mathbf{P}}\langle\mathbf{A}\times_{i}{\mathrm{P}}_{(i)}^{\perp},\mathbf{B}\times_{i}{\mathrm{P}}_{(i)}\rangle_{C}$ $=\langle{\mathrm{P}}_{(i)}^{\perp}A_{(i)},{\mathrm{P}}_{(i)}B_{(i)}\rangle_{\ell^{2}}$ (44)
@@
$\|\boldsymbol{\Phi}-\widehat{\boldsymbol{\Phi}}\|_{F}^{2}$ $\leq\|\boldsymbol{\Phi}\times_{1}{\mathrm{P}}_{(1)}^{\perp}\|_{F}^{2}+\|(\boldsymbol{\Phi}\times_{2}{\mathrm{P}}_{(2)}^{\perp})\times_{1}{\mathrm{P}}_{(1)}\|_{F}^{2}+\ldots$ (45)
@@
holds for a desired $\varepsilon\geq 0$. Such a bound is provided by the result in (38). However, in Algorithm 1 we can set $\varepsilon_{C}$ and $\varepsilon_{q}$, but not $\varepsilon_{s}$, which is implicitly determined by the the index set $\widetilde{\Omega}_{D}$. This motivates an adaptive slice sampling completion algorithm, which ensures that a proxy of (46) is fulfilled. For the adaptive completion we define a testing set of indexes ${\Omega}_{D}^{\rm test}\subset\Omega_{D}$ and gradually increase the training set $\widetilde{\Omega}_{D}$ up to the point when the completed tensor satisfies
@@
Output: $\rm{HTT}\big{(}\widetilde{\boldsymbol{\Phi}}\big{)}$ and the error $\|\boldsymbol{\Phi}|_{\Omega^{\rm test}}-\widetilde{\boldsymbol{\Phi}}|_{\Omega^{\rm test}}\|_{F}/\|\boldsymbol{\Phi}|_{\Omega^{\rm test}}\|_{F}$.
@@ 3.1 Projection TROM
For fast and accurate computations of trajectories $\mathbf{u}(t,\boldsymbol{\alpha})$ for any $\boldsymbol{\alpha}\in\mathcal{A}$, we consider a projection based ROM which uses the proposed HTT LRTD instead of the conventional POD as the dimension reduction technique. The HTT decomposition allows for the recovery of information about the parameter dependence of reduced spaces from a smaller set of pre-computed snapshots. This information is exploited for building parameter specific ROMs. The LRTD–ROM was introduced in mamonov2022interpolatory and further developed and analyzed in mamonov2024tensorial ; mamonov2023analysis . However, in those earlier works it is assumed that a full snapshot tensor $\boldsymbol{\Phi}$ is available and therefore its LRTD in one of the standard tensor rank revealing formats (e.g., CP, Tuker, and TT) can be computed.
For the projection TROM we aim to find a low dimensional subspace $V^{\ell}(\boldsymbol{\alpha})\in\mathbb{R}^{M_{1}}$ with $\ell=\dim(V^{\ell}(\boldsymbol{\alpha}))$, which is parameter-specific (e.g., varies depending on $\boldsymbol{\alpha}\in\mathcal{A}$) so that an approximation $\mathbf{u}^{\rm rom}(t,\boldsymbol{\alpha})$ to $\mathbf{u}(t,\boldsymbol{\alpha})$ is found by solving (1) projected onto $V^{\ell}(\boldsymbol{\alpha})$. This TROM solution $\mathbf{u}^{\rm rom}(t,\boldsymbol{\alpha})$ is then given by its coordinates in an orthogonal basis for $V^{\ell}(\boldsymbol{\alpha})$.
@@
In an interpolatory version of the TROM (see olshanskii2024approximating for a non-interpolatory version), one finds $V^{\ell}(\boldsymbol{\alpha})$ through interpolation procedure for the slices of $\widetilde{\boldsymbol{\Phi}}$. To define it, we first consider an interpolation operator that approximates a smooth function $g:\mathcal{A}\to\mathbb{R}$ using its values at the grid nodes ${\widehat{\mathcal{A}}}$. More precisely, we assume ${\boldsymbol{\chi}}^{i}\,:\,\mathcal{A}\to\mathbb{R}^{K_{i}},\quad i=1,\dots,D,$ such that for any continuous function $g:\mathcal{A}\to\mathbb{R}$,
@@
defines an interpolant for $g$. One straightforward choice is the Lagrange interpolation of order $p$: for any $\boldsymbol{\alpha}\in\mathcal{A}$, let $\widehat{\alpha}_{i}^{i_{1}},\ldots,\widehat{\alpha}_{i}^{i_{p}}$ be the $p$ closest grid nodes to $\alpha_{i}$ on $[\alpha_{i}^{\min},\alpha_{i}^{\max}]$, for $i=1,\ldots,D$. Then,
@@
If $\boldsymbol{\alpha}=\widehat{\boldsymbol{\alpha}}\in{\widehat{\mathcal{A}}}$, then ${\boldsymbol{\chi}}^{i}(\widehat{\boldsymbol{\alpha}})$ simply encodes the position of $\widehat{\alpha}_{i}$ among the grid nodes on $[\alpha^{\min}_{i},\alpha^{\max}_{i}]$. Therefore, for $\varepsilon=0$ the matrix $\widetilde{\Phi}(\widehat{\boldsymbol{\alpha}})$ is exactly the matrix of all snapshots for the particular $\widehat{\boldsymbol{\alpha}}$. For a general $\boldsymbol{\alpha}\in\mathcal{A}$, the matrix $\widetilde{\Phi}(\boldsymbol{\alpha})$ is the result of interpolation between snapshots (approximately) recovered by solving the completion problem. We will show that this interpolation is easy to do if $\widetilde{\boldsymbol{\Phi}}$ is in HTT format. Specifically, for an arbitrary given $\boldsymbol{\alpha}\in\mathcal{A}$ the parameter-specific local reduced space $V^{\ell}(\boldsymbol{\alpha})$ of dimension $\ell$ is the space spanned by the first $\ell$ left singular vectors of $\widetilde{\Phi}(\boldsymbol{\alpha})$:
@@ 3.2 In-tensor interpolation and finding the local basis
$[\mathrm{C}_{\chi}(\boldsymbol{\alpha})]_{i_{1},i_{2}}=\widetilde{\boldsymbol{\Phi}}_{\mathbf{i}}^{q}\times_{1}{\boldsymbol{\chi}}^{1}(\boldsymbol{\alpha})\dots\times_{D}{\boldsymbol{\chi}}^{D}(\boldsymbol{\alpha}),\quad\text{with}~{}\mathbf{i}=(i_{1},i_{2}).$ (54)
Practically, the tensor matrix products in (54) are computed as follows. Each $\mathbf{g}^{\mathbf{i}}_{j}$ from (26) is a $\widetilde{r}^{\mathbf{i}}_{j-1}\times K_{j}\times\widetilde{r}^{\mathbf{i}}_{j}$ tensor, then $G^{\mathbf{i}}_{j}=\mathbf{g}^{\mathbf{i}}_{j}\times_{2}{\boldsymbol{\chi}}^{j}(\boldsymbol{\alpha})$ is a $\widetilde{r}^{\mathbf{i}}_{j-1}\times\widetilde{r}^{\mathbf{i}}_{j}$ matrix and we calculate $[\mathrm{C}_{\chi}(\boldsymbol{\alpha})]_{i_{1},i_{2}}=G^{\mathbf{i}}_{1}\dots G^{\mathbf{i}}_{D}$ as a product of $D$ small size matrices.
@@
Note that the explicit computation of the basis $\mathbf{z}_{i}(\boldsymbol{\alpha})$ is not required for the online stage. Therefore, the essential information about $\widetilde{\boldsymbol{\Phi}}$ needed for the online part includes only the set of TT-tensors $\widetilde{\boldsymbol{\Phi}}_{\mathbf{i}}$:
@@ 3.3 Summary of the completion based TROM
    • Parameter grid ${\widehat{\mathcal{A}}}$ as in (4);
@@
      Execute Algorithm 2 with inputs ${\Omega}_{D}^{\rm test}$, $\widetilde{\Omega}_{D}$, $\varepsilon$, $N_{\max}$, and $P_{i}\in\mathbb{N}_{+}$, $i=1,\ldots,N_{\max}$, to compute slice sampling completion $\rm{HTT}\big{(}\widetilde{\boldsymbol{\Phi}}\big{)}$ using the numerical solver $\mathcal{S}$ in Step 4(c) for updating the data $\mathcal{D}$;
@@
      Project the system (1) onto the universal ROM space $\widetilde{U}$ using the orthonormal basis stored in $\widetilde{\mathrm{U}}^{(1)}$ from $\rm{HTT}\big{(}\widetilde{\boldsymbol{\Phi}}\big{)}$ to obtain the projected solver $\widetilde{\mathcal{S}}$ from the full order solver $\mathcal{S}$;
    Output: $\rm{HTT}\big{(}\widetilde{\boldsymbol{\Phi}}\big{)}$, projected solver $\widetilde{\mathcal{S}}$.
@@
    Input: $\mbox{\rm online}(\widetilde{\boldsymbol{\Phi}})$ as defined in (58), reduced space dimension $\ell\leq q_{\min}=\min\{q_{1},q_{2}\}$, parameter vector $\boldsymbol{\alpha}\in\mathcal{A}$, and projected solver $\widetilde{\mathcal{S}}$;
@@
      Use tensors $\widetilde{\boldsymbol{\Phi}}_{i}$ from $\mbox{\rm online}(\widetilde{\boldsymbol{\Phi}})$ to assemble the core matrix $\mathrm{C}_{\chi}(\boldsymbol{\alpha})\in\mathbb{R}^{q_{1}\times q_{2}}$ as in (54);

### v3

@@ 1 Introduction
we introduce a Cartesian grid on $\mathcal{A}$ by distributing $K_{i}$ nodes $\{\widehat{\alpha}_{i}^{j}\}_{j=1,\dots,K_{i}}$ within each of the intervals $\alpha_{i}^{\min},\alpha_{i}^{\max}]$ for $i=1,\dots,D$. The nodes of the grid form the set
@@
[…]
@@
It is clear that for a high dimension $D$ or a fine parameter grid $^{\widehat{\mathcal{A}}}$, working directly with $\boldsymbol{\Phi}$ is prohibitively expensive, if possible at all. Therefore, one approach to model order reduction for a system like (1) is based on the assumption that $\boldsymbol{\Phi}$ can be effectively approximated by a low-rank tensor $\widetilde{\boldsymbol{\Phi}}$ in one of the commonly used formats. Depending on the definition of tensor rank, various tensor decomposition formats can be employed, including CANDECOMP/PARAFAC (CP), Tucker (HOSVD), or Tensor Train (TT). While known rigorous analyses revealing how tensor ranks depend on the properties of a parameterized differential equation and the targeted accuracy of the solution manifold recovery, or its smoothness, are very limited (see, e.g., bachmayr2017kolmogorov and discussions in khoromskij2011tensor ; nouy2017low ; bachmayr2023low ), numerical evidence suggests that for many practical parametric PDEs, solutions to such problems or statistics derived from them are well approximated in low-rank formats kressner2011low ; ballani2015hierarchical ; dolgov2015polynomial ; ballani2016reduced ; eigel2017adaptive ; bachmayr2018parametric ; dolgov2019hybrid ; glau2020low ; mamonov2022interpolatory ; mamonov2024tensorial .
@@
with a prescribed $\varepsilon\geq 0$. The inexact completion (7) is a commonly used problem setup for the case of noisy data candes2010matrix and this is the formulation we are interested here. For CP, Tucker, and TT tensor ranks, the completion problems (6) and (7) are NP-hard. Building on the success of solving low-rank matrix completion problems candes2010power ; cai2010singular ; candes2012exact ; vandereycken2013low ; jain2013low ; nguyen2019low , a popular approach is to relax (6)–(7) into convex optimization problems signoretto2010nuclear ; gandy2011tensor ; bengua2017efficient . Other approaches to approximate tensor completion include ALS methods yokota2016smooth ; grasedyck2019stable , Riemannian optimization kressner2014low ; steinlechner2016riemannian , Bayesian methods zhao2015bayesian ; zhao2015bayesian2 , and projection methods rauhut2015tensor , with applications ranging from video recovery to seismic data reconstruction bengua2017efficient ; liu2022efficient .
@@
as a slice, hence the name slice sampling. For the problem (1), we have $C=2$, and in general we assume that $C$ is small so that using Cartesian grid in physical variables is feasible. The slice (8) coincides with the snapshot matrix (2), where $\boldsymbol{\alpha}=\big(\widehat{\alpha}_{1}^{k_{1}},\dots,\widehat{\alpha}_{D}^{k_{D}}\big)^{T}$. The sampled indices $k_{1},\dots,k_{D}$ form a sparse subset of the grid $^{\widehat{\mathcal{A}}}$.
@@
The reconstruction of scalar output quantities of parametric solutions in tensor format from incomplete observations was addressed in ballani2015hierarchical ; glau2020low . In ballani2015hierarchical the authors employed a tensor cross approximation, while glau2020low applied TT-completion via Riemannian optimization to recover an option pricing statistic from solutions of parametrized Heston and multi-dimensional Black-Scholes models. Additionally, a comparison of TT-cross interpolation and TT-completion for a parameterized diffusion equation in steinlechner2016riemannian demonstrated that TT-completion requires fewer PDE solver executions to find a low-rank approximation of a particular solution functional.
@@ 2 Slice sampling tensor completion
[…]
@@ 2.2 Tensor-train completion for partially sampled modes
$\boldsymbol{\Psi}\times_{n}\mathrm{M}]_{i_{1},\ldots,i_{n-1},j,i_{n+1},\ldots,i_{d}}=\sum_{k=1}^{m_{n}}\boldsymbol{\Psi}_{i_{1},\ldots,i_{n-1},k,i_{n+1},\ldots,i_{d}}\mathrm{M}_{jk},\quad j=1,\ldots,m.$ (19)
@@
[…]
@@ 2.5 Accuracy of HTT completion
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\langle\mathbf{A}\times_{i}{\mathrm{P}}_{(i)}^{\perp},\mathbf{B}\times_{i}{\mathrm{P}}_{(i)}\rangle_{F}}$ $=\langle{\mathrm{P}}_{(i)}^{\perp}A_{(i)},{\mathrm{P}}_{(i)}B_{(i)}\rangle_{\ell^{2}}$ (44)
@@
$\|\boldsymbol{\Phi}-\widehat{\boldsymbol{\Phi}}\|_{F}^{2}$ $=\|\boldsymbol{\Phi}\times_{1}{\mathrm{P}}_{(1)}^{\perp}\|_{F}^{2}+\|(\boldsymbol{\Phi}\times_{2}{\mathrm{P}}_{(2)}^{\perp})\times_{1}{\mathrm{P}}_{(1)}\|_{F}^{2}+\ldots$ (45)
@@
holds for a desired $\varepsilon\geq 0$. Such a bound is provided by the result in (38). In Algorithm 1 we can set $\varepsilon_{C}$ and have a control of $\varepsilon_{q}$ depending on the fitting algorithm used for pointwise tensor completion. However $\varepsilon_{s}$ is implicitly determined by the the index set $\widetilde{\Omega}_{D}$. This motivates an adaptive slice sampling completion algorithm, which ensures that a proxy of (46) is fulfilled. For the adaptive completion we define a testing set of indexes ${\Omega}_{D}^{\rm test}\subset\Omega_{D}$ and gradually increase the training set $\widetilde{\Omega}_{D}$ up to the point when the completed tensor satisfies
@@
[…]
@@ 3.1 Projection TROM
For fast and accurate computations of trajectories $\mathbf{u}(t,\boldsymbol{\alpha})$ solving (1) for any $\boldsymbol{\alpha}\in\mathcal{A}$, we consider a projection based ROM which uses the proposed HTT LRTD instead of the conventional POD as the dimension reduction technique. The HTT decomposition allows for the recovery of information about the parameter dependence of reduced spaces from a smaller set of pre-computed snapshots. This information is exploited for building parameter specific ROMs. The LRTD–ROM was introduced in mamonov2022interpolatory and further developed and analyzed in mamonov2024tensorial ; mamonov2023analysis . However, in those earlier works it is assumed that a full snapshot tensor $\boldsymbol{\Phi}$ is available and therefore its LRTD in one of the standard tensor rank revealing formats (e.g., CP, Tuker, and TT) can be computed.
For the sake of exposition, we assume that all spatial degrees of freedom are vectorized into a single mode, so that only space and time remain as physical modes; i.e., $C=2$ for the rest of this section. In projection TROM we aim to find a low dimensional subspace $V^{\ell}(\boldsymbol{\alpha})\in\mathbb{R}^{M_{1}}$ with $\ell=\dim(V^{\ell}(\boldsymbol{\alpha}))$, which is parameter-specific and an approximation $\mathbf{u}^{\rm rom}(t,\boldsymbol{\alpha})$ to $\mathbf{u}(t,\boldsymbol{\alpha})$ is found by solving (1) projected onto $V^{\ell}(\boldsymbol{\alpha})$. This TROM solution $\mathbf{u}^{\rm rom}(t,\boldsymbol{\alpha})$ is then given by its coordinates in an orthogonal basis for $V^{\ell}(\boldsymbol{\alpha})$.
@@
In an interpolatory version of the TROM (see olshanskii2024approximating for a non-interpolatory version), one finds $V^{\ell}(\boldsymbol{\alpha})$ through interpolation procedure for the slices of $\widetilde{\boldsymbol{\Phi}}$. To define it, we first consider an interpolation operator that approximates a smooth function $g:\mathcal{A}\to\mathbb{R}$ using its values at the grid nodes $^{\widehat{\mathcal{A}}}$. More precisely, we assume ${\boldsymbol{\chi}}^{i}\,:\,\mathcal{A}\to\mathbb{R}^{K_{i}},\quad i=1,\dots,D,$ such that for any continuous function $g:\mathcal{A}\to\mathbb{R}$,
@@
defines an interpolant for $g$. One straightforward choice is the Lagrange interpolation of order $p$: for any $\boldsymbol{\alpha}\in\mathcal{A}$, let $\widehat{\alpha}_{i}^{i_{1}},\ldots,\widehat{\alpha}_{i}^{i_{p}}$ be the $p$ closest grid nodes to $\alpha_{i}$ on $\alpha_{i}^{\min},\alpha_{i}^{\max}]$, for $i=1,\ldots,D$. Then,
@@
[…]
@@ 3.2 In-tensor interpolation and finding the local basis
$\mathrm{C}_{\chi}(\boldsymbol{\alpha})]_{i_{1},i_{2}}=\widetilde{\boldsymbol{\Phi}}_{\mathbf{i}}^{q}\times_{1}{\boldsymbol{\chi}}^{1}(\boldsymbol{\alpha})\dots\times_{D}{\boldsymbol{\chi}}^{D}(\boldsymbol{\alpha}),\quad\text{with}~\mathbf{i}=(i_{1},i_{2}).$ (54)
Practically, the tensor matrix products in (54) are computed as follows. Each $\mathbf{g}^{\mathbf{i}}_{j}$ from (26) is a $\widetilde{r}^{\mathbf{i}}_{j-1}\times K_{j}\times\widetilde{r}^{\mathbf{i}}_{j}$ tensor, then $G^{\mathbf{i}}_{j}=\mathbf{g}^{\mathbf{i}}_{j}\times_{2}{\boldsymbol{\chi}}^{j}(\boldsymbol{\alpha})$ is a $\widetilde{r}^{\mathbf{i}}_{j-1}\times\widetilde{r}^{\mathbf{i}}_{j}$ matrix and we calculate $\mathrm{C}_{\chi}(\boldsymbol{\alpha})]_{i_{1},i_{2}}=G^{\mathbf{i}}_{1}\dots G^{\mathbf{i}}_{D}$ as a product of $D$ small size matrices.
@@
Note that the explicit computation of the basis functions $\mathbf{z}_{i}(\boldsymbol{\alpha})$ is not required during the online stage. In particular, the projection of the dynamical system onto the space $V^{\ell}(\boldsymbol{\alpha})=\mathrm{span}\{\mathbf{z}_{i}(\boldsymbol{\alpha})\}_{i=1}^{\ell}$ is performed using only the matrix of basis coordinates $\widetilde{\mathrm{U}}^{(1)}$; see Algorithm 3. Therefore, the essential information about $\widetilde{\boldsymbol{\Phi}}$ needed for the online part includes only the set of TT-tensors $\widetilde{\boldsymbol{\Phi}}_{\mathbf{i}}$:
@@ 3.3 Summary of the completion based TROM
    • Parameter grid $^{\widehat{\mathcal{A}}}$ as in (4);
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
3.4 Complexity of CTROM
The computational requirements of the completion-based TROM are determined by the following steps:
  • (i)
    Offline stage: Completion of $\widetilde{\boldsymbol{\Phi}}$ in the HTT format;
  • (ii)
[…]
  • (iii)
    Online stage: Computation of the coordinates of the parameter-specific reduced basis for a given $\boldsymbol{\alpha}$;
  • (iv)
    Solution of (1) projected onto the reduced space.
Since step (iv) is common to all projection-based ROMs, we focus on steps (i)–(iii). The offline stage is dominated by the completion algorithm and repeated FOM solves used to generate the training set. The number of FOM calls scales at least as $O(\sum_{j=1}^{D}r^{\rm max}_{j}K_{j}r^{\rm max}_{j+1})$, where $r^{\rm max}_{j}=\max_{\mathbf{i}\in\Omega_{C}^{q}}r^{\mathbf{i}}_{j}$. The cost of the completion step depends on the chosen scheme; in the numerical experiments we employ the stable rank-adaptive ALS (SALSA) method from grasedyck2019stable , with complexity $O(\sum_{j=1}^{D}|\Omega_{C}^{q}||\widetilde{\Omega}_{D}||r^{\rm max}_{j}r^{\rm max}_{j+1}|^{2})$ per adaptive step.
[…]
In step (iii), the cost of computing the $\boldsymbol{\alpha}$-specific reduced basis is governed by the interpolation procedure and the evaluation of the leading $n$ left singular vectors of the core matrix. Since the vectors $\mathbf{e}^{i}(\boldsymbol{\alpha})$ are sparse (typically $p=2$ or $3$ nonzero entries in Cartesian sampling), forming the core matrix $\mathrm{C}(\boldsymbol{\alpha})$ requires $O\Big(\sum_{i=2}^{D}\widetilde{r}_{i-1}\widetilde{r}_{i}\widetilde{r}_{i+1}\Big)$ operations. The subsequent SVD involves only the small $\widetilde{r}_{1}\times\widetilde{r}_{D+1}$ matrix. If a reduced basis in the physical space is required, the vectors are recovered as linear combinations of the columns of $\mathrm{U}$, with cost $O(|\Omega_{C}^{q}|\widetilde{r}_{1}\ell)$.